# RankMixer: Scaling Up Ranking Models in Industrial Recommenders

### v2

@@ Abstract.
Recent progress on large language models (LLMs) has spurred interest in scaling up recommendation systems, yet two practical obstacles remain. First, training and serving cost on industrial Recommenders must respect strict latency bounds and high QPS demands. Second, most human-designed feature-crossing modules in ranking models were inherited from the CPU era and fail to exploit modern GPUs, resulting in low Model Flops Utilization (MFU) and poor scalability. We introduce RankMixer, a hardware-aware model design tailored towards a unified and scalable feature-interaction architecture. RankMixer retains the transformer’s high parallelism while replacing quadratic self-attention with multi-head token mixing module for higher efficiency. Besides, RankMixer maintains both the modeling for distinct feature subspaces and cross-feature-space interactions with Per-token FFNs. We further extend it to one billion parameters with a Sparse-MoE variant for higher ROI. A dynamic routing strategy is adapted to address the inadequacy and imbalance of experts training. Experiments show RankMixer’s superior scaling abilities on a trillion-scale production dataset. By replacing previously diverse handcrafted low-MFU modules with RankMixer, we boost the model MFU from 4.5% to 45%, and scale our online ranking model parameters by 70x while maintaining roughly the same inference latency. We verify RankMixer’s universality with online A/B tests across two core application scenarios (Recommendation and Advertisement). Finally, we launch 1B Dense-Parameters RankMixer for full traffic serving without increasing the serving cost, which improves user active days by 0.3% and total in-app usage duration by 1.08%.
@@ 1. Introduction
  • Leveraging levers of high MFU and performance-optimization, we scale the model parameters by 100× without increasing the inference cost, including improving GPU utilization, SparseMoE acceleration, and low-precision inference.
  • We conduct extensive offline and online experiments and investigate the model’s scaling law on the trillion-scale industrial recommendation dataset. The RankMixer model has been successfully deployed on the Douyin Feed Recommendation Ranking for full-traffic serving, achieving 0.2% and 0.5% increase on active days and App duration, respectively.
@@ 3.3.1. Multi-head Token Mixing
Although self-attention has proven highly effective in large language models, we find it to be suboptimal for recommendation systems. In self-attention, attention weights are calculated using the inner product of tokens. This method works well for NLP, where all tokens share a unified embedding space. However, in recommendation tasks, the feature spaces are inherently heterogeneous. Computing an inner-product similarity between two heterogeneous semantic spaces is notoriously difficult—particularly in recommender systems, where the ID space of the features from user and item side may contain hundreds of millions of elements. Consequently, applying self-attention to such diverse inputs does not outperform the parameter-free multi-head Token Mixing approach and consumes more computations and IO operations.
@@ 3.3.2. Per-token FFN
It is noticed that after the multi-head Token Mixing, each token consists of exactly one head of all original tokens. It is because the original tokens are mapped into different subspaces that the tokens contain different information after mixing. We observe that tokens after mixing tend to collapse into one same representation if a parameter-shared FFN is applied. Besides, per-token FFN enhances the modeling ability by introducing more parameters while keeping the computational complexity unchanged.
@@ 3.4. Sparse MoE in RankMixer
To further increase the scaling ROI, we can replace the dense FFNs of each pertoken with Sparse Mixture-of-Experts (MoE) blocks, so that the model’s capacity grows while computation cost remains roughly constant. Vanilla Sparse-MoE, however, degrades in RankMixer because: (i) uniform $k$-expert routing. Top-$k$ selection treats all feature tokens equally, wasting the budget on low-information tokens and starving high-information ones, which hinders the model from capturing differences between tokens. (ii) expert under-training. Per-token FFNs already multiply parameters by #tokens; adding non-shared experts further explodes the expert count, producing highly unbalanced routing and poorly trained experts;
@@ 4.3. Scaling Laws of different models
To be Sepcific, We can scale up RankMixer model by increasing width ($D$), feature token ($T$) and Layer($L$). In our experiments we observed a conclusion shared with LLM scaling laws: model quality correlates primarily with the total number of parameters, and different scaling directions (depth L, width D, tokens T) yield almost identical performance. From a compute-efficiency standpoint, while larger hidden-dim generate larger matrix-multiply shapes and thus achieve higher MFU than stacking more layers. So the final configuration for 100M and 1B are set as ($D$=768, $T$=16, $L$=2) and ($D$=1536, $T$=32, $L$=2) respectively.
@@ 4.4. Ablation Study
   w/o skip connections $-0.07\%$
   w/o multi-head token mixing $-0.50\%$
   w/o layer normalization $-0.05\%$
   Per-token FFN $\rightarrow$ shared FFN $-0.31\%$
@@ 4.5. Sparse-MoE Scalability and Expert Balance
Table 4. Online AB test result lasting 5 months for Feed Recommendation Scenarios on different activeness user groups in both Douyin and Douyin lite app
@@
Low-active +0.457% +0.8594% +0.6564% +1.7523% +0.951% +0.4248% +2.1946% +1.327% +3.2622% +0.6988%
Middle-active +0.4318% +1.1855% +0.6778% +1.9561% +0.972% +0.4115% +1.8366% +1.738% +2.31% +1.0317%
High-active +0.1239% +0.4917% +0.272% +1.3125% +0.37% +0.0666% +0.8427% +2.187% +1.556% +2.5246%
Overall +0.2008% +0.4996% +0.2862% +1.6016% +0.3827% +0.1648% +0.726% +0.8401% +1.319% +0.8269%
@@
Metric Base-16M RankMixer-1B Change
# Dense Param 15.8M 1.1B $\uparrow$ 70$\times$
@@
Inference Precisioin fp32 fp16 $\uparrow$ 2$\times$
@@ 4.6. Online Serving cost
How can we prevent inference latency from scaling linearly with a 100 $\times$ increase in parameters? In practical systems, latency is inversely proportional to throughput and directly proportional to additional machine resources. We take advantaget of two key levers in the model-system co-design to decouple the parameter Growth from inference cost:
@@
    FLOPs-to-parameter ratio. The fourth column of Table 6 reports the number of floating-point operations (FLOPs) required per parameter. Thanks to the per-token FFN design, RankMixer expands parameter capacity without proportionally increasing computation, achieving only one-third the FLOPs/parameter ratio of the baseline—a 3 $\times$ efficiency gain.
@@
    Model FLOPs Utilization (MFU). Also shown in Table 6, MFU indicates how close the kernels run to the GPU’s peak throughput. By using large GEMM shapes and good parallel topology, RankMixer raises MFU by nearly 10 $\times$, shifting the model from an Memory-bound to a Compute-bound regime.
Combining these levers, a parameter growth of over 100 $\times$ translates into only a 3 $\times$ increase in serving latency. Besides, we uses further engineering optmization to make the deployed 1-B parameter RankMixer to run faster than the previous production baseline including:
  • Pertoken FFN Op fusion: Multiple parallel per-token FFNs can be merged into a three-dimensional tensor, allowing several per-token FFNs to be fused into a single computational operator to reduce the launch overhead of kernels. This strategy yields a 30% speed-up.
  • Mixed-precision inference: The primary computations in RankMixer consist of several large matrix multiplications that are well suited for half-precision (fp16) inference as mentioned before, while operations sensitive to precision, such as layer normalization, continue to use full precision. This optmization increase the throughput by 45% and reduce the latency by 31.5%
  • Acceleration on SparseMoE: By implementing custom sparse-GEMM kernels, we converted the per-token FFN into a $1/8$-sparsity Sparse-MoE, cutting end-to-end inference latency by more than 40%.
@@ 4.7. Online Performance
The Previous baselines in these scenarios are 10 Million-level parameter DLRM model. We replaced the Dense part with the RankMixer-1B model. We conduct online A/B test experiments across Feed Recommendation and Advertising. The A/B test result on Feed Recommendation are shown on the Table 4 and the result of Advertising is shown on the Table 5.
RankMixer was deployed and evaluated in the three personalised-ranking application including the Feed Recommendation (RankMixer-1B) and Advertising (RankMixer-1B). RankMixer delivered statistically significant uplifts on all business-critical metrics. We can also observe that the improvements on low-active users are largest on the Table 4 compared to other activeness level user groups with over 0.45% Active Days improvements, which demonstrates the model’s strong generalization capability. These results demonstrate that the RankMixer as a unified backbone generalises reliably to different application across all important scenarios.
@@ 5. Conclusion
In this article, we introduce our latest RankMixer model that has been fully deployed on Douyin Feed ranking. It combines the model designs for heterogeneous feature interactions and the highly parallelizable architecture for serving efficiency. Experiments have demonstrated its superior performance and steep scaling law. After its full deployment on the Douyin app, it has obtained 0.2% and 0.5% increase on active days and App duartion.

### v3

@@ Abstract.
Recent progress on large language models (LLMs) has spurred interest in scaling up recommendation systems, yet two practical obstacles remain. First, training and serving cost on industrial Recommenders must respect strict latency bounds and high QPS demands. Second, most human-designed feature-crossing modules in ranking models were inherited from the CPU era and fail to exploit modern GPUs, resulting in low Model Flops Utilization (MFU) and poor scalability. We introduce RankMixer, a hardware-aware model design tailored towards a unified and scalable feature-interaction architecture. RankMixer retains the transformer’s high parallelism while replacing quadratic self-attention with multi-head token mixing module for higher efficiency. Besides, RankMixer maintains both the modeling for distinct feature subspaces and cross-feature-space interactions with Per-token FFNs. We further extend it to one billion parameters with a Sparse-MoE variant for higher ROI. A dynamic routing strategy is adapted to address the inadequacy and imbalance of experts training. Experiments show RankMixer’s superior scaling abilities on a trillion-scale production dataset. By replacing previously diverse handcrafted low-MFU modules with RankMixer, we boost the model MFU from 4.5% to 45%, and scale our online ranking model parameters by two orders of magnitude while maintaining roughly the same inference latency. We verify RankMixer’s universality with online A/B tests across two core application scenarios (Recommendation and Advertisement). Finally, we launch 1B Dense-Parameters RankMixer for full traffic serving without increasing the serving cost, which improves user active days by 0.3% and total in-app usage duration by 1.08%.
@@ 1. Introduction
  • Leveraging levers of high MFU and performance-optimization, we scale the model parameters by 70× without increasing the inference cost, including improving MFU and Quantization.
  • We conduct extensive offline and online experiments and investigate the model’s scaling law on the trillion-scale industrial recommendation dataset. The RankMixer model has been successfully deployed on the Douyin Feed Recommendation Ranking for full-traffic serving, achieving 0.3% and 1.08% increase on active days and App duration, respectively.
@@ 3.3.1. Multi-head Token Mixing
Although self-attention has proven highly effective in large language models, we find it to be suboptimal for recommendation systems. In self-attention, attention weights are calculated using the inner product of tokens. This method works well for NLP, where all tokens share a unified embedding space. However, in recommendation tasks, the feature spaces are inherently heterogeneous. Computing an inner-product similarity between two heterogeneous semantic spaces is notoriously difficult—particularly in recommender systems, where the ID space of the features from user and item side may contain hundreds of millions of elements. Consequently, applying self-attention to such diverse inputs does not outperform the parameter-free multi-head Token Mixing approach and consumes more computations, Memory IO operations and GPU memory usage.
@@ 3.3.2. Per-token FFN
Compared to the parameter-all-shared FFN, per-token FFN enhances the modeling ability by introducing more parameters while keeping the computational complexity unchanged.
It is worth emphasizing that a Per-token FFN differs from MMoE experts in that each Per-token FFN sees a distinct token input, whereas all experts in a MMoE share the same input. Unlike a MMoE, where many experts process the same input, and unlike a Transformer, where different input share one FFN, RankMixer splits the inputs and the parameters simultaneously which is good for learning diversity in different feature sub-spaces.
@@ 3.4. Sparse MoE in RankMixer
[…]
@@ 4.3. Scaling Laws of different models
To be Sepcific, We can scale up RankMixer model by increasing width ($D$), feature token ($T$) and $L$). In our experiments we observed a conclusion shared with LLM scaling laws: model quality correlates primarily with the total number of parameters, and different scaling directions (depth L, width D, tokens T) yield almost identical performance. From a compute-efficiency standpoint, while larger hidden-dim generate larger matrix-multiply shapes and thus achieve higher MFU than stacking more layers. So the final configuration for 100M and 1B are set as ($D$=768, $T$=16, $L$=2) and ($D$=1536, $T$=32, $L$=2) respectively.
@@ 4.4. Ablation Study
[…]
@@ 4.5. Sparse-MoE Scalability and Expert Balance
Table 4. Online AB test result for Feed Recommendation Scenarios in both Douyin and Douyin lite app and the improvements are all statistically significant. According to long-term reverse A/B tests and continuous observations over long-term reverse experiments, the gains provided by RankMixer-1B have not yet converged, and the results presented here are still improving.
@@
Overall +0.2908% +1.0836% +2.3852% +1.9874% +0.7886% +0.1968% +0.9869% +1.1318% +2.0744% +1.1338%
Low-active +1.7412% +3.6434% +8.1641% +4.5393% +2.9368% +0.5389% +2.1831% +4.4486% +3.3958% +0.9595%
Middle-active +0.7081% +1.5269% +2.5823% +2.5062% +1.2266% +0.4235% +1.9011% +1.7491% +2.6568% +0.6782%
High-active +0.1445% +0.6259% +1.828% +1.4939% +0.4151% +0.0929% +1.1356% +1.8212% +1.7683% +2.3683%
@@
Metric OnlineBase-16M RankMixer-1B Change
#Param 15.8M 1.1B $\uparrow$ 70$\times$
@@
Hardware FLOPs fp32 fp16 $\uparrow$ 2$\times$
@@ 4.6. Online Serving cost
How can we prevent inference latency from exploding with a two-order-of-magnitude increase in parameters? In practical systems, latency is inversely proportional to throughput and directly proportional to the cost of serving machine resources. Compared to our previous fully-deployed 16M-parameter model (with a structure integrated with DLRM and DCN), our RankMixer model scaled parameters by approximately 70$\times$ to 1B. Despite this significant parameter increase, the final inference latency remained stable due to our hardware-aligned model design and optimization strategies.
When greatly increasing model parameters, latency can be decomposed into the following formula:
$\text{Latency}=\frac{\#\text{Param}\times\text{FLOPs/Param ratio}}{\text{MFU}\times\text{(Theoretical Hardware FLOPs)}}$
As illustrated in Table 6, the two-order-of-magnitude parameter increase is progressively counteracted by 3.6$\times$ decrease of FLOPs/Param ratio, 10 $\times$ increase of MFU and Quatization-based 2 $\times$ Hardware FlOPs improvements.
@@
    FLOPs/Param ratio. The three row of Table 6 reports the number of floating-point operations (FLOPs) required per parameter. Thanks to the model design, RankMixer achieves a 70-fold increase in parameters with only around a 20-fold increase in FLOPs, achieving only one-third the FLOPs/Param ratio of the baseline—a 3.6,$\times$ efficiency gain. In other words, RankMixer can accommodate three times more parameters than the baseline at the same FLOPs budget.
@@
    Model FLOPs Utilization (MFU). Also shown in Table 6, MFU indicates the utilization of machine computing. By using large GEMM shapes, good parallel topology (fusing parallel per-token FFNs into one kernel), reduce the memory bandwith cost and overhead, RankMixer raises MFU by nearly 10 $\times$, shifting the model from an Memory-bound to a Compute-bound regime.
  3. (3)
    Quantization Half-precision (fp16) inference will increase the theoretical peak Hardware FLOPS of GPUs by 2 $\times$.The primary computations in RankMixer consist of several large matrix multiplications that are well suited for half-precision (fp16) inference as mentioned before
@@ 4.7. Online Performance
The Previous baselines in these scenarios are 16M parameter model combined with DLRM and DCN. We replaced the Dense part with the RankMixer-1B model improving a 0.7% AUC. We conduct online A/B test experiments across Feed Recommendation and Advertising. The long-term observation of A/B test result on Feed Recommendation for 8 months are shown on the Table 4  [^1] This result was updated on July 24, 2025, since the long-term experiments indicating the gains had not yet saturated. and the result of Advertising is shown on the Table 5.
RankMixer was deployed and evaluated in the three personalised-ranking application including the Feed Recommendation (RankMixer-1B) and Advertising (RankMixer-1B). RankMixer delivered statistically significant uplifts on all business-critical metrics. We can also observe that the improvements on low-active users are largest on the Table 4 compared to other activeness level user groups with over 1.7412% Active Days improvements, which demonstrates the model’s strong generalization capability. These results demonstrate that the RankMixer as a unified backbone generalises reliably to different application scenarios.
@@ 5. Conclusion
In this article, we introduce our latest RankMixer model that has been fully deployed on Douyin Feed ranking. It combines the model designs for heterogeneous feature interactions and the highly parallelizable architecture for serving efficiency. Experiments have demonstrated its superior performance and steep scaling law. After its full deployment on the Douyin app, it has obtained 0.3% and 1% increase on active days and App duartion.